OpenStudy (anonymous):

Given the function f(x)=2−4x−x^2, calculate f(3+h)−f(3)/h

OpenStudy (anonymous):

\[f(3)=2-12-9=-19\] \[f(3+h)=2-4(3+h)-(3+h)^2=-h^2-10 h-19\] \[f(3+h)-f(3)=-h^2-10h\]

OpenStudy (anonymous):

divide by h, get \[-h-10\]